OpenStudy (anonymous):

how would i go about solving 8e^2x=20

OpenStudy (anonymous):

natural log cancels the e. Then solve normally.

OpenStudy (anonymous):

divide by 8, take the log, divide by 2

OpenStudy (anonymous):

Like 8/20 first. Then ln e^2x, and you get 2x=ln(20/8). Then you divide all tht by 2 and that's the answer.

OpenStudy (anonymous):

I mean 20/8

OpenStudy (anonymous):

\[8e ^{2x} = 20\] \[\ln (20/8) = 2x\] \[ x= .5*\ln(20/8) = \ln(\sqrt{2.5})\]

OpenStudy (anonymous):

thank youuuuuuuu